Antiques R Us is a mature manufacturing firm. The company just paid a dividend of n $11, but management expects to reduce the payout by 8 percent per year indefinitely. If you require a return of 15 percent on this stock, what will you pay for a share today? Multiple Choice $47.83 $43.56 $44.00 $44.44 $144.57
Answer #1

Current price=D1/(Required return-Growth rate)

=(11*(1-0.08))/(0.15-(0.08)

=10.12/0.23

=$44
Answer #2

Current price=D1/(Required return-Growth rate)

=(11*(1-0.08))/(0.15+(0.08)

=10.12/0.23

=$44